Consider the following two projects

Project 1: High rise residential building project

Project 2: Low rize building

First Cost, $

320,000

540,000

M&O Cost, $/year

45,000

35,000

Benefits, $/year

110,000

150,000

Disbenefits, $/year

20,000

45,000

Life, years

10

20

      -       A.       B.       C.       D.       E.       F.       G.       H.   

For project 1, the Benefits/Cost ratio is nearest to

      -       A.       B.       C.       D.       E.       F.       G.       H.   

For project 2, the Benefits/Cost ratio is nearest to

      -       A.       B.       C.       D.       E.       F.       G.       H.   

Which project must be chosen?

A.

0.90

B.

0.85

C.

1.20

D.

1.15

E.

1.10

F.

Project 1

G.

Project 2

H.

0.95
Answer #1

It is essential to understand the meaning of each term before we perform any calculation.

First cost can be taken as a fixed cost.Whether the project is in operation for one year,or ten years,this cost is initially incurred.

M&O cost is the management and operations cost per year.Say addition of new machinery,maintenance cost or cost of depreciation.

Benefits include the revenue, increase in productivity of the employees,the benefits of the buyers,etc.

Dis-benefits may mean negative externalities,etc.

Now,the cost benefit analysis is done for the entire duration of the project.i.e.10 years for project 1 and 20 years for Project 2.But comparing the benefits and costs over different life spans may cause discrepancies.We thus construct project 1 over two years to facilitate comparison.Thus multiplying all calculations of project 1 by 2.Initial cost is also assumed to be incurred twice for project 1(since the life of Project1 is two years)

Total cost = Initial cost(First Cost)+M&O cost

Net Benefits=Benefits-Disbenefits

For Project 1

Total cost=[3,20,000+(45,000*10)]*2=7,70,000*2

Net benefits=[(110000*10)-(20000*10)]*2=9,00,000*2

Benefit-Cost Ratio=(9,00,000*2)/(7,70,000*2)=1.168 or 1.15(nearest to D)

For Project 2,

Total cost=540000+(35000*20)=12,40,000

Net benefits=(150000*20)-(45000*20)=21,00,000

Benefit Cost Ratio=21,00,000/1,240,000=1.693 or 1.20(although far,but nearest to 1.20,i.e. Option C)

Because Project 2 has a higher Benefit -Cost Ratio,it is the preferred project. Option G is chosen.